Circular No. 6798

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971227 P. M. Woods, University of Alabama in Huntsville; C. Kouveliotou,

Universities Space Research Association; and G. J. Fishman, Marshall Space Flight Center, NASA, report on behalf of the BATSE Team: “GRB 971227 (IAUC 6796, 6797) was recorded as BATSE trigger No. 6546. It was a weak, short event of duration ∼ 6 s that exhibits two well distinguished peaks, separated by 2.5 s. The burst fluence (above 25 keV) was 9.3 (± 1.4) × 10^{−7} erg cm^{−2} and its peak flux (between 50 and 300 keV) was 3.3 ± 0.2 photons cm^{−2} s^{−1}. The burst was not detected above 300 keV.”

SUPERNOVA 1997ef IN UGC 4107 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics; and D. Balam, University of Victoria, report: “Recent CCD images of SN 1997ef obtained with the Whipple Observatory 1.2-m telescope by P. Berlind, with the MDM 2.4-m telescope by J. Thorstensen, and with the Dominion Astrophysical Observatory’s 1.8-m Plaskett telescope indicate that this object has begun to decline in brightness. Visual magnitude estimates relative to the star (V = 17.05) located 1′ southwest of UGC 4107: Dec. 16.3 UT, 16.56; 27.3, 16.84; 28.3, 16.94. Spectra obtained at the Tillinghast telescope by Berlind and L. Macri, and at the MDM 2.4-m telescope by Thorstensen, between Dec. 17 and 27 show that the rapid velocity shift seen in the major spectral features has slowed. After correcting for the redshift of UGC 4107, a number of peaks can now be identified at 397 nm (Ca II), 457 nm (?), 486 nm (Hβ), 547 nm (?), 587 nm (He I), 614 nm (?), and 655 nm (Hα). The Hα emission peak is rounded, while the unidentified maxima have a triangular shape. Currently, the spectrum appears similar to the peculiar type-Ib supernovae 1988L (Filippenko 1988, A.J. 96, 1941) and 1983I (Wheeler et al. 1987, Ap.J. 313, L69), which also displayed broad maxima near the wavelength of Hα. These objects and SN 1997ef appear to be extreme examples of massive stars that have lost all but a small amount of hydrogen through mass loss.”

COMET C/1997 D1 (MUELLER) Visual m_{1} estimates: Nov. 1.00 UT, 11.7 (M. Lehký, Hradec Králové,

Czech Republic, 0.42-m reflector); Dec. 21.83, 12.1 (Lehký); 28.66, 13.4 (A. Pearce, Cadoux, W. Australia, 0.41-m reflector).

1997 December 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green